# Experiential Explanations for Reinforcement Learning

### v4

@@ Abstract
Reinforcement Learning (RL) systems can be complex and non-interpretable, making it challenging for non-AI experts to understand or intervene in their decisions. This is due in part to the sequential nature of RL in which actions are chosen because of future rewards. However, RL agents discard the qualitative features of their training, making it difficult to recover user-understandable information for “why” an action is chosen. We propose a technique Experiential Explanations to generate counterfactual explanations by training influence predictors along with the RL policy. Influence predictors are models that learn how sources of reward affect the agent in different states, thus restoring information about how the policy reflects the environment. A human evaluation study revealed that participants presented with experiential explanations were better able to correctly guess what an agent would do than those presented with other standard types of explanation. Participants also found that experiential explanations are more understandable, satisfying, complete, useful, and accurate. Qualitative analysis provides insights into the factors of experiential explanations that are most useful. [^1] Code repository available at:
https://github.com/amal994/Experiential-Explanations-RL
@@ 1 Introduction
Reinforcement Learning (RL) techniques are becoming increasingly popular in critical domains such as robotics and autonomous vehicles. However, applications such as autonomous cars and socially assistive robots in healthcare and home settings are expected to interact with non-AI experts. People often seek explanations for abnormal events or observations to improve their understanding of the agent and better predict and control its actions [1]. This need for an explanation will likely arise when interacting with these systems, whether it is a car taking an unexpected turn or a robot performing a task unconventionally. Without an explanation, users can find it difficult to trust the agent’s ability to act safely and reasonably  [2], especially if they have limited artificial intelligence expertise. This can even be in the case where the agent is operating optimally and without failure; the agent’s optimal behavior may not match the user’s expectations, resulting in confusion and lack of trust. Therefore, RL systems need to provide good explanations without compromising their performance. To address this need, explainable RL has become a rapidly growing area of research with many open challenges [3, 4, 5, 6, 7, 8, 9].
But what makes a good explanation for end-users who cannot change the underlying model? How do we generate explanations if we cannot modify the underlying RL model or degrade its performance? In RL, an agent learns dynamically to maximize its reward through a series of experiences interacting with the environment [10]. Through these interactions, the agent builds up a model of utility, $Q(s,a)$, which estimates the future reward that can be expected to be achieved if a particular action $a$ is performed from a particular state $s$.
While an RL agent learns from its experiences during training, those experiences are inaccessible after their training is over [4]. This is because $Q(s,a)$ summarizes future experiences as a single, real-valued number; this is all that is needed to execute a policy $\pi=\mathrm{argmax}_{a}Q(s,a)$. For example, consider a robot that receives a negative reward for being close to the stairs and thus learns that states along an alternative route have higher utility. When it comes time to figure out why the agent executed one action trajectory over another, the policy is devoid of any information from which to construct an explanation beyond the fact that some actions have lower expected utility than others.
@@
We propose a technique, Experiential Explanations, in which deep RL agents generate explanations in response to on-demand, local, counterfactuals proposed by users. These explanations qualitatively contrast the agent’s intended action trajectory with a trajectory proposed by the user, and specifically link agent behavior to environmental contexts. The agent maintains a set of models—called influence predictors—of how different sparse reward states influence the agent’s state-action utilities and, thus, its policy. Influence predictors are trained alongside an RL policy with a specific focus on retaining details about the training experience. The influence models are “outside the black box”, but provide information with which to interpret the agent’s black-box policy in a human-understandable fashion.
Consider the illustrative example in Figure 1 where the user observes the agent go down around the wall. The user expects the agent to go up, but the agent goes down instead. The user asks why the agent did not choose up, which appears to lead to a shorter route to the goal. One possible and correct explanation would be that the agent’s estimated expected utility for the down action is higher than that for the up action. However, this is not information that a user can easily act upon. The alternative, experiential explanation, states that up will pass through a region that is in proximity to dangerous locations. The user can update their understanding that the agent prefers to avoid stairs. The user can also understand how to take an explicit action to change the environment, such as blocking stairs. Our technique focuses on post-hoc, real-time explanations of agents operating in situ, instead of pre-explanation of agent plans [11], although the two are closely related.
We evaluate our Experiential Explanations technique with studies with human participants. We show that explanations generated by our technique allow users to better understand and predict agent actions and are found to be more useful. We additionally perform a qualitative analysis of participant responses to see how users use the explanations in our technique and baseline alternatives to reason about the agent’s behavior, which not only provides evidence for Experimental Explanations, but provides a blueprint to understand the human factors of other explanation techniques.
@@ 2 Background and Related work
Reinforcement learning is an approach to learning in which an agent attempts to maximize its reward through feedback from a trial-and-error process. RL is suitable for Markov Decision Processes, which can be represented as a tuple $M=\langle S,A,T,R,\gamma\rangle$ where $S$ is the set of possible world states, $A$ is the set of possible actions, $T$ is the state transition function, which determines the probability of the following state $P(S)$ as a function of the current state and action. $T:S\times A\rightarrow P(S)$, $R$ is the reward function $R:S\times A\rightarrow R$, and $\gamma$ is a discount factor $0\leq\gamma\leq 1$. RL first learns a policy $\pi:S\rightarrow A$, which defines which actions should be taken in each state. Deep RL uses deep neural networks to estimate the expected future utility such that $\pi(s)=\operatorname*{argmax}_{a}Q_{\theta}(s,a)$ where $\theta$ are the model parameters.
@@
As deep reinforcement learning is increasingly used in sensitive areas such as healthcare and robotics, Explainable RL (XRL) is becoming a crucial research field. For an overview of XRL, please refer to Milani et al. [8]. We highlight the most relevant work on XRL to our approach.
Rationale generation [12] is an approach to explanation generation for sequential environments in which explanations approximate how humans would explain what they would do in the same situation as the agent. Rationale generation does not “open the black box” but looks at the agent’s state and action and applies a model trained from human explanations. Rationales do not guarantee faithfulness. However, our models train alongside the agent and achieve a high degree of faithfulness.
@@
Some XRL techniques generate explanations by decomposing utilities and showing information about the components of the utility score for an agent’s state [13, 14, 15, 16]. Our technique is related to decompositions in that we learn the influence of different sources of reward on utilities, avoiding the trade-off between using a semi-interpretable system or a better black-box system.
@@
Our work focuses on counterfactual explanations, but we differ in the explanation goals and approach from other works. For example, Frost et al. [17] trained an exploration policy to generate test time trajectories to explain how the agent behaves in unseen states after training. Explanation trajectories help users understand how the agent would perform under new conditions. While this method gives users a global understanding of the agent, our method focuses on providing local counterfactual explanations of an agent’s action. Contrastive approaches such as van der Waa et al. [18] leverage an interpretable model and a learned transition model. Madumal et al. [19] explain local actions with a causal structural model as an action influence graph to encode the causal relations between variables. Sreedharan et al. [20] generate contrastive explanations using a partial symbolic model approximation, including knowledge about concepts, preconditions, and action costs. These methods provided answers to user’s local why not questions. Our approach minimizes the use of predefined tasks or a priori environment knowledge.
Olson et al. [21] explains the local decisions of an agent by demonstrating a counterfactual state in which the agent takes a different action to illustrate the minimal alteration necessary for a different result. Huber et al.[22] furthers this concept by creating counterfactual states with an adversarial learning model. Our technique, however, concentrates on the qualitative distinctions between trajectories and not on the states.
@@ 3 Experiential Explanations
Experiential Explanations are contrastive explanations generated for local decisions with minimal reliance on structured inputs and without imposing limitations on the agent or the RL algorithm. Our explanation technique uses additional models called influence predictors that learn how sparse rewards affect the agent’s utility predictions. In essence, influence predictors tell us how strongly or weakly any source of reward (positive or negative) is impacting states in the state space that the agent believes it will pass through. This is in contrast to the agent’s learned policy, which aggregates the utilities with respect to all rewards. This enables agents that use our explanation technique to provide additional context to the choices they make because it knows about their utility landscape in finer-grained detail. As in Figure 1, we see the explanation referencing the agent’s relationship to the environment in terms of negative and positive elements.
Once we have trained these influence predictor models, explanation generation proceeds in two phases. First, given a decision made by the agent and a request for an explanation, we generate a state-action trajectory that the agent will take to maximize the expected reward plus a counterfactual trajectory based on the user’s expectation. For example, in Figure 1 the agent’s trajectory is down, but the user asks about going up.
Second, we use influence predictors to compare the two trajectories and extract information about the different influences along each. In the following sections, we will dive into the details about the influence predictors, what they are, how we train them, how to use them to generate explanations, and the other components of the explanation generation system and how we generate explanations.
@@ 3.1 Influence Predictors
Influence predictors reconstruct the effects of the received rewards on utility during training. These models are trained alongside the agent to predict the strength of the influence of different sources of reward on the agent: $U_{c}(s,a)$ where each $c\in C$ is a distinct source (or class) of reward. For example, a terminal goal state might be a source of positive reward, stairs might be a source of negative reward, and other dangerous objects may be other sources of negative reward. In a more complex environment, multiple sources of positive and negative rewards can be helpful to outline the agent’s plan. For instance, a robot attempting to make a cup of coffee would receive positive rewards for obtaining a clean mug, hot milk and coffee, then combining the milk and coffee in the mug and stirring them. Negative influences could include getting a dirty cup or adding salt.
In principle, any agent architecture and optimization algorithm can be used with our explanation technique as long as we can observe its transitions during training. Influence predictors are trained similarly to an RL agent (e.g., standard Deep $Q$-networks (DQN) [23], or A2C [24]) but with two differences. First, during training, each influence predictor model can only receive rewards from one class of rewards. Influence predictors are trained using loss based on a modified Bellman equation: $U_{c}(s_{t},a_{t})=|r^{c}_{t+1}|+\gamma\max_{a^{\prime}}U_{c}(s_{t+1},a^{\prime})$ where $r^{c}_{t+1}$ is the reward associated with $c$ delivered to the agent upon transition to state $s_{t+1}$. The absolute value re-interprets the reward as an influence, since the influence predictor is now predicting the utility of attaining the reward, even if it is negative. As a consequence of ever receiving only the absolute value of a class of rewards, $U_{c}(\cdot)$ is an estimate of the strength of $c$ in $s$ and $a$.
Second, influence predictors are not used to guide the agent’s exploration during training or inference. They rely on the main policy’s training loop to generate transitions.
@@
Each source of reward has its own influence predictor model. Typically, one class is associated with a goal state (or states required for task completion) and produces positive rewards. Sometimes an intermediate positive reward is given when it is known in advance that the task requires particular actions or states to be traversed, as in the case of reward shaping [25]. Other classes are associated with states that are detrimental to the task and produce negative rewards.
@@
The training process for influence predictors is shown in Algorithm  1. The agent, controlled by its base policy learning algorithm, interacts with the environment, executing actions and receiving state observations. State transitions, actions, and rewards are recorded and used to populate separate replay buffers for each influence predictor. Each time step, we sample each replay buffer and backpropagate loss in each influence model.
@@
             $transitions_{T}[t][reward]\leftarrow\lvert transitions_{T}[t][reward]\rvert$
@@
             $transitions_{T}[t][reward]\leftarrow 0$
@@
To illustrate, Figure 2 shows an agent operating in an environment with two reward classes: a positive reward class for the terminal goal, and a negative reward class for the blue squares. The influence map on the left was generated by querying the negative influence predictor for each state and maxing across all possible actions. The negative influence trends are stronger closer to the blue squares. The influence map in the middle is generated from the positive influence predictor. The map on the right is the agent’s main policy model, trained on all reward classes as normal. Even though the visualization aggregates action utilities, one can see that the agent will prefer a trajectory in the upper part of the map.
@@ 3.2 Explanation Generation
At any time during the execution of the policy, the user can request a counterfactual explanation by proposing an action $(a_{\mathrm{U}})$ rather than the one the agent chose $(a_{\mathrm{A}})$. Given the action chosen by the agent and the alternative action proposed by the user, the explanation generation process proceeds in three phases.
@@
First, we generate and record the trajectory $\tau_{\mathrm{A}}$ the agent will take from this point forward. [^2] We assume access to a simulator or a world model (e.g., [26]) though a world-model based learner was not used to create a minimalistic experiment. Second, we generate a counterfactual trajectory $\tau_{\mathrm{U}}$ as if the agent had executed the user’s proposed alternative action. To generate $\tau_{\mathrm{U}}$, we force the agent to follow the user’s suggested alternative action (or sequence of actions in the case that the user wants to specify a partial or complete counterfactual trajectory) and complete the trajectory using the same planning approach.
@@
Finally, we use influence predictors to compare the two trajectories. Recall, in Section 3.1, we have an influence predictor $IP_{c}$ per reward source $c\in C$. We have access to all the influence predictor values in both trajectories at this stage, which gives us control over the types of explanation we can generate depending on our use case. We can produce local, detailed, global, or aggregated explanations. Furthermore, explanations can be enriched by giving names to reward states and textual descriptions of the consequences of visiting a reward state (e.g., “stairs” and “fall down”, respectively). In this work, we explore two variants of the explanations: an aggregated explanation and a detailed local explanation.
@@
Aggregated Explanation. An aggregated explanation presents an overview of the effect of each reward class on the agent throughout a complete trajectory. To generate these explanations, we calculate the mean value of each influence predictor $IP_{c}$ for all states $s\in\tau_{\mathrm{A}}$ and compare this mean with the other trajectories
@@
$\operatorname*{argmax}_{a\in\{a_{\mathrm{U}},a_{\mathrm{A}}\}}(\frac{\sum_{s}{\max(IP_{c}(\tau_{\mathrm{a}}[s]))}}{length(\tau_{\mathrm{a}})}),\forall c\in C.$
@@
The resulting explanation assigns each influence predictor to the trajectory in which they were the most dominant. We then take these values and translate them into English explanations. If the values generated by $IP_{c}$ were similar for both trajectories, then they can be excluded from the explanation, as they do not provide valuable information. For example, in the case of two influence predictors in a simple environment, if we assume that we only got a difference in the negative influence predictor. The explanation would be as follows: “If I go up, I fear falling down the stairs; going down feels safer”. This case is demonstrated in Figure 1. Suppose that there is a difference in an additional $IP_{c}$ mean. In that case, we add it to the explanation by appending its relevant sentence to explain all influence predictors if we observe a difference in their mean values.
@@
Local Explanation. In more complex environments where there are many influences, the explanations can shed light on the agent’s near-horizon and concentrate more on exploring local values. These local explanations zoom in on the most influential factors related to actual and counterfactual actions over their corresponding segments, which are denoted by $t_{\mathrm{A}}\subseteq\tau_{\mathrm{A}}$. Each segment starts at the beginning of the designated action and includes a pre-determinted number of additional steps (we used five steps) after the end of the action to capture some of the longer-term effects.
@@
$\{\operatorname*{argmax}_{c\in\{C\}}(\max{(IP_{c}(s))})|\forall a\in\{a_{\mathrm{U}},a_{\mathrm{A}}\},\forall s\in t_{\mathrm{a}}\}$
From the result, we obtain a set that includes all unique influence predictors $IP_{c}$ that had a maximum value at any point of the segment. If there is no difference between the segments of $a_{\mathrm{A}}$ and $a_{\mathrm{U}}$ with respect to their sets of maximum influences $IP_{c}$, we resolve to compare the mean values of the influence predictors during the same period and obtain the three influences with the highest means in their segments. First, we get the mean of every influence predictor in each segment:
$mean(IP_{c},a)=\frac{\sum_{s}{\max{(IP_{c}(t_{\mathrm{A}}[s]))}}}{||t_{\mathrm{A}}||},\forall c\in C,\forall a\in\{a_{U},a_{\mathrm{A}}\}$
Then, we obtain the top three influences for each segment $t_{\mathrm{A}}$ and $t_{\mathrm{U}}$, ordered by $mean(IP_{c})$.
We provide this explanation even if there is no difference in the top influences, which can happen if the alternative action considered is too similar to the actual action taken. To generate English explanations, we utilize the list of top influence predictors obtained through either the maximum or highest mean method. An example of a generated explanation can be: “If I got the nearby cup, I would still have to clean it before making my coffee. But if I get a cup from the cupboard, I will not need to clean it and I can immediately make coffee.”

### v5

@@ Abstract
Reinforcement Learning (RL) systems can be complex and non-interpretable, making it challenging for non-AI experts to understand or intervene in their decisions. This is due in part to the sequential nature of RL in which actions are chosen because of their likelihood of obtaining future rewards. However, RL agents discard the qualitative features of their training, making it difficult to recover user-understandable information for “why” an action is chosen. We propose a technique Experiential Explanations to generate counterfactual explanations by training influence predictors along with the RL policy. Influence predictors are models that learn how different sources of reward affect the agent in different states, thus restoring information about how the policy reflects the environment. Two human evaluation studies revealed that participants presented with Experiential Explanations were better able to correctly guess what an agent would do than those presented with other standard types of explanation. Participants also found that Experiential Explanations are more understandable, satisfying, complete, useful, and accurate. Qualitative analysis provides information on the factors of Experiential Explanations that are most useful and the desired characteristics that participants seek from the explanations. [^1] Data availability statement: The authors declare that the code and data supporting the findings of this study are available within the paper and its supplementary information files.
@@ 1 Introduction
Reinforcement Learning (RL) techniques are becoming increasingly popular in critical domains such as robotics and autonomous vehicles. However, applications such as autonomous cars and socially assistive robots in healthcare and home settings are expected to interact with non-AI experts. People often seek explanations for abnormal events or observations to improve their understanding of the agent and better predict and control its actions [1]. This need for an explanation will likely arise when interacting with these systems, whether it is a car taking an unexpected turn or a robot performing a task unconventionally. Without an explanation, users can find it difficult to trust the agent’s ability to act safely and reasonably  [2], especially if they have limited artificial intelligence expertise. Even in the case where the agent is operating optimally and without failure, the agent’s optimal behavior may not match the user’s expectations, resulting in confusion and lack of trust. Therefore, RL systems need to provide good explanations without compromising their performance. To address this need, explainable RL has become a rapidly growing area of research with many open challenges [3, 4, 5, 6, 7, 8, 9].
But what makes a good explanation for end-users who cannot change the underlying model? How do we generate explanations if we cannot modify the underlying RL model or degrade its performance? In RL, an agent learns dynamically to maximize its reward through a series of experiences interacting with the environment [10]. Through these interactions, the agent builds a utility model, $V(s,a)$, which estimates the future reward that can be expected to be achieved if a particular action $a$ is performed from a particular state $s$.
While an RL agent learns from its experiences during training, those experiences are inaccessible after their training is over [4]. This is because $V(s,a)$ summarizes future experiences as a single, real-valued number; this is all that is needed to execute a policy $\pi=\mathrm{argmax}_{a}V(s,a)$. For example, consider a robot that receives a negative reward for being close to the stairs and thus learns that states along an alternative route have higher utility. When it comes time to figure out why the agent executed one action trajectory over another, the policy is devoid of any information from which to construct an explanation beyond the fact that some actions have lower expected utility than others.
@@
We propose a technique, Experiential Explanations, in which deep RL agents generate explanations in response to on-demand, local, counterfactuals proposed by users. These explanations qualitatively contrast the agent’s intended action trajectory with a trajectory proposed by the user and specifically link agent behavior to environmental contexts. The agent maintains a set of models—called influence predictors—of how different sparse reward states influence the agent’s state-action utilities and, thus, its policy. Influence predictors are trained alongside an RL policy with a specific focus on retaining details about the training experience. The influence models are “outside the black box”, but provide information with which to interpret the agent’s black-box policy in a human-understandable fashion.
Consider the illustrative example in Figure 1 where the user observes the agent go down around the wall. The user expects the agent to go up, but the agent goes down instead. The user asks why the agent did not choose up, which appears to lead to a shorter route to the goal. One possible and correct explanation would be that the agent’s estimated expected utility for the down action is higher than that for the up action. However, this is not information that a user can easily act upon. The alternative, Experiential Explanation, states that up will pass through a region that is in proximity to dangerous locations. The user can update their understanding that the agent prefers to avoid stairs. The user can also understand how to take an explicit action to change the environment, such as blocking stairs—an action a non-expert can take to influence agent behavior when they do not have the ability to directly change the agent. Our technique focuses on post-hoc, real-time explanations of agents operating in situ, instead of pre-explanation of agent plans [11], although the two are closely related.
We evaluate our Experiential Explanations technique with studies with human participants. We show that explanations generated by our technique allow users to better understand and predict agent actions and are found to be more useful. We additionally perform a qualitative analysis of participant responses to elicit how users use the information in the explanations in our technique and baseline alternatives to reason about the agent’s behavior. This not only provides evidence for Experimental Explanations, but provides a blueprint to understand the human factors of other explanation techniques.
@@ 2 Background and Related work
Reinforcement learning is an approach to learning in which an agent attempts to maximize its reward through feedback from a trial-and-error process. RL is suitable for Markov Decision Processes, which can be represented as a tuple $M=\langle S,A,T,R,\gamma\rangle$ where $S$ is the set of possible world states, $A$ is the set of possible actions, $T$ is the state transition function, which determines the probability of the following state $P(S)$ as a function of the current state and action. $T:S\times A\rightarrow P(S)$, $R$ is the reward function $R:S\times A\rightarrow\mathbb{R}$, and $\gamma$ is a discount factor $0\leq\gamma\leq 1$. RL first learns a policy $\pi:S\rightarrow A$, which defines which actions should be taken in each state. Deep RL uses deep neural networks to estimate the expected future utility such that $\pi(s)=\operatorname*{argmax}_{a}V_{\theta}(s,a)$ where $\theta$ are the model parameters.
@@
As deep reinforcement learning is increasingly used in sensitive areas such as healthcare and robotics, Explainable RL (XRL) is becoming a crucial research field. However, the reinforcement learning paradigm differs substantially from other machine learning paradigms, such as supervised learning, in that they solve sequential decision-making problems where each decision an agent makes is not an end to itself but establishes conditions for future actions; understanding any one decision may be dependent on understanding prior decisions as well as future potential decisions. For an overview of XRL, please refer to Milani et al. [8]. We highlight the most relevant work on XRL to our approach.
@@
Explanation via Reward Decomposition breaks down rewards into a sum of semantically meaningful reward categories, allowing actions to be evaluated based on the trade-offs between these categories [12]. Many XRL techniques generate explanations by decomposing utilities in the agent’s network and showing information about the components of the utility score for an agent’s state [13, 14, 15]. Our technique is related to decompositions in that we learn the influence of different sources of reward on utilities. However, we differ in that we approximate the decomposition using external models, avoiding the trade-off between using a semi-interpretable system or a better black-box system.
@@
Explanations are inherently contrastive, as they are framed in the context of particular counterfactual comparisons. Rather than merely questioning why event P took place, individuals ask why event P occurred rather than event Q [1]. Contrastive approaches such as van der Waa et al. [16] leverage an interpretable model and a learned transition model. Madumal et al. [17] explain local actions with a causal structural model as an action influence graph to encode the causal relations between variables. Sreedharan et al. [18] generate contrastive explanations using a partial symbolic model approximation, including knowledge about concepts, preconditions, and action costs. These methods provided answers to user’s local why not questions. While our approach still produces counterfactual explanations, we minimize the use of predefined symbolic or a priori environment knowledge. Other work looked at utilizing counterfactual explanations for other explanation goals. For example, Frost et al. [19] trained an exploration policy to generate test time trajectories to explain how the agent behaves in unseen states after training. Explanation trajectories help users understand how the agent would perform under new conditions. While this method gives users a global understanding of the agent, our method focuses on providing local counterfactual explanations of an agent’s action. Olson et al. [20] explains the local decisions of an agent by demonstrating a counterfactual state in which the agent takes a different action to illustrate the minimal alteration necessary for a different result. Huber et al.[21] furthers this concept by creating counterfactual states with an adversarial learning model. Our technique, however, concentrates on the qualitative distinctions between trajectories and not on the states.
Rationale generation, introduced by Ehsan et al. [22], is an approach to explanation generation for sequential environments in which explanations approximate how humans would explain what they would do in the same situation as the agent. Rationale generation does not “open the black box” but looks at the agent’s state and action and applies a model trained from human explanations. Rationales do not guarantee faithfulness, though studies show they can still provide actionable insights. Like rationale generation, our models train alongside the agent. However, our technique also achieve a high degree of faithfulness. State2Explanation (S2E), introduced in Das et al.[23], demonstrates generating more faithful explanations by using them to inform the agent’s reward shaping and requires the derivation of the concepts from mathematical representations and expert domain knowledge. By aligning state-action pairs with high-level concepts, S2E uses natural language templates to provide actionable insights during training and deployment. In our approach, we avoid intervening with the learning of the agent while preserving a high level of faithfulness by confining the learning only from the agent’s transitions and trying to minimize the reliance on domain knowledge to only the names and effects of reward components.
@@ 3 Experiential Explanations
Experiential Explanations are contrastive explanations generated for local decisions with minimal reliance on structured inputs and without imposing limitations on the agent or the RL algorithm. Our explanation technique uses additional models called influence predictors that learn how the sources of the sparse rewards affect the agent’s utility predictions. The agent’s learned policy learns the utilities with respect to all rewards. Influence predictors can tell us how strongly or weakly any source of reward (positive or negative) is impacting states in the state space that the agent believes it will pass through. The influence predictors provide the additional context that explains the agent’s choices through finer-grained utility detail. As in Figure 1, we see the explanation referencing the agent’s relationship to the environment in terms of negative and positive elements.
The main components of Experiential Explanations are the trained influence predictors and the two modules of explanation generation. The first module starts with the user proposing an alternative action and asking why the agent did something else; seeking a counterfactual explanation. This triggers simulation of the trajectories of both actions. For example, in Figure 1 the agent’s action is down, but the user asks about going up. The generated trajectories account for those actions and continue until the end of the respective episodes. In the second module, we compare the states along a specified horizon of the simulated trajectories using influence predictors.
In Section 3.1 we dive into the deeper details of the influence predictors. In Section 3.2 we demonstrate how they can be used in the explanation generation modules to generate explanations.
@@ 3.1 Influence Predictors
Influence predictors reconstruct the effects of rewards received on the utility learned by the agent during training. These models are trained alongside the agent to predict the strength of the influence of different sources of reward on the agent by learning a value function that outputs the expected utility from each source of reward: $U_{c}(s,a)$ where each $c\in C$ is a distinct source (or class) of reward. For example, a terminal goal state might be a source of positive reward, stairs might be a source of negative reward, and other dangerous objects may be other sources of negative reward. In a more complex environment, multiple sources of positive and negative rewards can be helpful in describing the agent’s plan. For instance, a robot attempting to make a cup of coffee would receive positive rewards for obtaining a clean mug, hot milk, and coffee, then combining the milk and coffee in the mug and stirring them. Negative influences could include getting a dirty cup or adding salt.
@@
Each source of reward has its own influence predictor model. Typically, one class is associated with a goal state (or states required for task completion) and produces positive rewards. Sometimes an intermediate positive reward is given when it is known in advance that the task requires particular actions or states to be traversed, as in the case of reward shaping [24]. Other classes are associated with states that are a priori known to be detrimental to the task and produce negative rewards.
In principle, Influence Predictors can be used with any agent architecture and optimization algorithm as long as we can observe its transitions during training. Influence predictors are trained similarly to an RL agent (e.g., standard Deep $Q$-networks (DQN) [25], or A2C [26]) but with two differences. First, influence predictors are not used to guide the agent’s exploration during training or inference. They rely on the main policy’s training loop to generate transitions. Second, each influence predictor model can receive only rewards from one class of rewards. Influence predictors are trained using loss based on a modified Bellman equation:
$U_{c}(s_{t},a_{t})=|r^{c}_{t+1}|+\gamma\max_{a^{\prime}}U_{c}(s_{t+1},a^{\prime})$ (1)
where $r^{c}_{t+1}$ is the reward associated with $c$ delivered to the agent upon transition to state $s_{t+1}$. The absolute value re-interprets the reward as an influence, since the influence predictor is now predicting the utility of attaining the reward, even if it is negative. As a consequence of ever receiving only the absolute value of a class of rewards, $U_{c}(\cdot)$ is an estimate of the strength of $c$ in $s$ and $a$.
@@
The training process for influence predictors is shown in Algorithm  1. The agent, controlled by its base policy learning algorithm, interacts with the environment, executing actions, and receiving state observations. State transitions, actions, and rewards are recorded and used to populate separate buffers for each influence predictor. For model updates, we sample the transitions from each designated buffer separately and backpropagate loss in each influence predictor.
@@
             $transition_{t}[c][reward]\leftarrow\lvert transitions_{T}[t][reward]$|$\rvert$|
@@
             $transition_{t}[c][reward]\leftarrow 0$
@@
To illustrate, Figure 2 shows an agent operating in an environment with two reward classes: a positive reward class for the terminal goal, and a negative reward class for the blue squares—hence, two influence predictors. The influence map on the left was generated by querying the negative influence predictor for each state and maxing across all possible actions. The negative influence trends are stronger closer to the blue squares. The influence map in the middle is generated from the positive influence predictor. The map on the right is the agent’s main policy model, trained on all reward classes as normal. Even though the visualization aggregates action utilities, one can see that the agent will prefer a trajectory in the upper part of the map.
@@ 3.2 Explanation Generation
At any time during the execution of the policy, the user can request a counterfactual explanation by proposing an action $(a_{\mathrm{user}})$ rather than the one the agent chose $(a_{\mathrm{agent}})$. Given the action chosen by the agent and the alternative action proposed by the user, the explanation generation process proceeds in three phases.
@@
First, we generate and record the trajectory $\tau_{\mathrm{agent}}$ the agent will take from this point forward. [^2] We assume access to a simulator or a world model (e.g., [27]) with sufficient fidelity to predict action outcomes, though a world-model based learner was not used to create a minimalistic experiment. Second, we generate a counterfactual trajectory $\tau_{\mathrm{user}}$ as if the agent had executed the user’s proposed alternative action. To generate $\tau_{\mathrm{user}}$, we force the agent to follow the user’s suggested alternative action (or sequence of actions in the case that the user wants to specify a partial or complete counterfactual trajectory) and complete the trajectory using the same planning approach.
@@
Finally, we use influence predictors to compare the two trajectories. Recall, in Section 3.1, we have an influence predictor $IP_{c}$ per reward source $c\in C$. We have access to all the influence predictor values in both trajectories at this stage, which gives us control over the types of explanation we can generate depending on our use case.
We can produce local, detailed, global, or aggregated explanations in natural language. To do so we require one modification to the typical Markov Decision Process formulation: we use an enhanced reward function,
$R:S\times A\rightarrow\mathbb{R}\times\mathcal{L}^{*}$ (2)
where $\mathcal{L}$ is a natural language vocabulary and $\mathcal{L}^{*}$ is the set of all possible natural language strings. The enchanced reward function produces both a real valued number and a natural language string that describes the circumstances under which the reward is given. For example, for the scenario in Figure 1, the enhanced reward function might produce the value $-1$ and “fall down stairs” when the agent encounters the location in the top center of the environment.
These textual descriptions of reward circumstances can be embedded in the explanation templates to automatically add stance or other details about the reward influences to the explanations. In this work, we explore two variants of the explanations: an aggregated explanation and a detailed local explanation.
@@
Aggregated Explanation. An aggregated explanation presents an overview of the effect of each reward class on the agent throughout a complete trajectory. To generate these explanations, we calculate the mean value of each influence predictor $IP_{c}$ for all states $s\in\tau_{\mathrm{agent}}$ and compare this mean with the other trajectories.
@@
$\operatorname*{argmax}_{a\in\{a_{\mathrm{user}},a_{\mathrm{agent}}\}}(\frac{\sum_{s}{\max(IP_{c}(\tau_{\mathrm{a}}[s]))}}{length(\tau_{\mathrm{a}})}),\forall c\in C.$ (3)
@@
The resulting explanation assigns each influence predictor to the trajectory in which they were the most dominant. We then take these values and translate them into natural language (English) explanations. If the values generated by $IP_{c}$ were similar for both trajectories, then they can be excluded from the explanation, as they do not provide valuable, contrastive information. For example, in the case of two influence predictors in a simple environment, if we assume that we only had a difference in the negative influence predictor, the explanation would be as follows: “If I go up, I fear falling down the stairs; going down feels safer”. This case is demonstrated in Figure 1. Suppose that there is a difference in an additional $IP_{c}$ mean. In that case, we add it to the explanation by appending its relevant sentence to explain all influence predictors if we observe a significant difference in their mean values.
@@
Local Explanation. In more complex environments where there are many influences, the explanations can shed light on the agent’s near-horizon and concentrate more on exploring local values. These local explanations zoom in on the most influential factors related to actual and counterfactual actions over their corresponding segments, which are denoted by $t_{\mathrm{agent}}\subseteq\tau_{\mathrm{agent}}$. Each segment starts at the beginning of the designated action and includes a pre-determinted number of additional steps (we used five steps) after the end of the action to capture some of the longer-term effects.
@@
$\{\operatorname*{argmax}_{c\in\{C\}}(\max{(IP_{c}(s))})|\forall a\in\{a_{\mathrm{user}},a_{\mathrm{agent}}\},\forall s\in t_{\mathrm{agent}}\}$ (4)
From equation 4 we get two sets of influence predictors that had the maximum value at any state within the segments $t_{agent}$ and $t_{user}$. If there is no difference between the segments of $a_{\mathrm{agent}}$ and $a_{\mathrm{user}}$ with respect to their sets of maximum influences $IP_{c}$, we resolve to compare the mean values of the influence predictors during the same period and obtain the three influences with the highest means in their segments. First, we get the mean of every influence predictor in each segment:
$mean(IP_{c},a)=\frac{\sum_{s}{\max{(IP_{c}(t_{\mathrm{agent}}[s]))}}}{||t_{\mathrm{agent}}||},\forall c\in C,\forall a\in\{a_{\mathrm{user}},a_{\mathrm{agent}}\}$ (5)
Then, we obtain the top three influences for each segment $t_{\mathrm{agent}}$ and $t_{\mathrm{user}}$, ordered by $mean(IP_{c})$.
We provide this explanation even if there is no difference in the top influences, which can happen if the alternative action considered is too similar to the actual action taken. To generate English explanations, we use the list of the top influence predictors obtained through the maximum- or highest-mean-method. And then embed them into the templates. An example of a generated explanation can be: “If I got the nearby cup, I would still have to clean it before making my coffee. But if I get a cup from the cupboard, I will not need to clean it and I can immediately make coffee.”. Where “cleaning” and “making coffee” are terms derived from the enhanced reward sources, as described above.